# A Comparative Analysis of Wealth Index Predictions in Africa between three Multi-Source Inference Modelsthanks: Please cite the ECMLPKDD’24 version of this paper

### v1

@@ 0.A.1 Performance on test sets
Table A1: Reported validation R2 scores for each approach. M1 reports validation results on census data, as shown in Fig. S4 of the supplementary information [2]. M2 presents validation results for single-country training in Table 1 [16]. We retrain the M3 models for each country using a subset of the original features reported in [8], and report the performance (R2) of the best model, indicated by the lowest root-mean-square error, RMSE (see details in Section 2.2).
Country / R2 M1 M2 M3
Sierra Leone 83.0 89.82 80.7
Liberia - 89.3 76.3
Uganda - 89.02 78.5
Rwanda 82.0 87.41 67.5
South Africa - 73.03 49.5
Gabon - - 81.9
0.A.2 Overall predictions
(a) High-resolution
(b) Districts
(c) IWI distribution
Figure A1: IWI predictions in Sierra Leone (M3: CB). We expect M1 and M2 to produce wealthier maps than M3  (Table 3), with M1 being the wealthiest (Table 2). Wealth predictions shown for M1, are linearly rescaled. Color bars display the average predicted IWI scores, centered at the mean of M3’s predictions for comparison. (a) Prediction at the original scale. (b) Aggregated mean wealth at the administrative level 2. Compared to M3, M2 is consistently predicting lower wealth, while M1 predicts higher values (M1 and M3 fulfill the expectation). (c) Predicted wealth distribution (From top: PDF, CDF). IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [22]) is similar between M2 and M3, but differs for M1.
(a) High-resolution
(b) Districts
(c) IWI distribution
Figure A2: IWI predictions in Liberia (M3: CB). We expect both M1 and M2 to produce slightly wealthier maps than M3  (Table 3). Wealth predictions shown for M1, are linearly rescaled. Color bars display the average predicted IWI scores, centered at the mean of M3’s predictions for comparison. (a) Prediction at the original scale. (b) Aggregated mean wealth at the administrative level 2. Compared to M3, M2 is consistently predicting lower wealth, while M1 predicts higher values (M1 and M3 fulfill the expectation). (c) Predicted wealth distribution (From top: PDF, CDF). IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [22]) differs across models.
(a) High-resolution
(b) Districts
(c) IWI distribution
Figure A3: IWI predictions in Uganda (M3: CBw). We expect M3 to produce slightly wealthier maps than M1 and comparable to M2 (Table 3). Wealth predictions shown for M1, are linearly rescaled. Color bars display the average predicted IWI scores, centered at the mean of M3’s predictions for comparison. (a) Prediction at the original scale. (b) Aggregated mean wealth at the administrative level 2. Compared to M3, M2 is consistently predicting lower wealth, while M1 predicts slightly lower values (M1 and M3 fulfill the expectation). (c) Predicted wealth distribution (From top: PDF, CDF). IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [22]) is similar between M1 and M3, but differs for M2.
(a) High-resolution
(b) Districts
(c) IWI distribution
Figure A4: IWI predictions in Rwanda (M3: CBw). We expect both M1 and M2 to produce slightly poorer poverty maps than M3 (Table 3). Wealth predictions shown for M1, are linearly rescaled. Color bars display the average predicted IWI scores, centered at the mean of M3’s predictions for comparison. (a) Prediction at the original scale. (b) Aggregated mean wealth at the administrative level 2. Compared to M3, both M2 and M1 predict lower wealth values. However, M2 is twice as poor than M1 (M1 and M3 fulfill the expectation). (c) Predicted wealth distribution (From top: PDF, CDF). IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [22]) is similar between M1 and M3, but differs for M2.
(a) High-resolution
(b) Districts
(c) IWI distribution
Figure A5: IWI predictions in Gabon (M3: CNNa+CB). We expect M1’s poverty map to be wealthier than M3 (Table 3). Wealth predictions shown for M1, are linearly rescaled. Color bars display the average predicted IWI scores, centered at the mean of M3’s predictions for comparison. (a) Prediction at the original scale. (b) Aggregated mean wealth at the administrative level 2. Compared to M3, M1 predicts a wealther country (as expected). (c) Predicted wealth distribution (From top: PDF, CDF). IWI poverty line ($35^{th}$ percentile, corresponding to poverty headcount ratios at \$1.25 a day [22]) is similar between M1 and M3, but differs for M2.
0.A.3 Overlapping locations
(a) M1 vs. M2
(b) M1 vs. M3
(c) M2 vs. M3
Figure A6: IWI predictions in overlapping locations in Sierra Leone (M3: CB). Left: Correlation between the predictions of two models. Right: Overlapping locations. Colors indicate differences in predicted values per model pair. M1 generally overestimates wealth compared to M2 (blue in (a)) and M3 (blue in (b)), and predictions are more homogeneous. In (c), M2’s predictions are more heterogeneous than M3’s, despite achieving relatively low RMSE.
(a) M1 vs. M2
(b) M1 vs. M3
(c) M2 vs. M3
Figure A7: IWI predictions in overlapping locations in Liberia (M3: CB). Left: Correlation between the predictions of two models. Right: Overlapping locations. Colors indicate differences in predicted values per model pair. M1 generally overestimates wealth compared to M2 (blue in (a)) and M3 (blue in (b)), and predictions are more homogeneous. In (c), M2’s predictions are more heterogeneous than M3’s.
(a) M1 vs. M2
(b) M1 vs. M3
(c) M2 vs. M3
Figure A8: IWI predictions in overlapping locations in Uganda (M3: CBw). Left: Correlation between the predictions of two models. Right: Overlapping locations. Colors indicate differences in predicted values per model pair. M1 generally overestimates wealth compared to M2 (blue in (a)) and M3 (blue in (b)). Predictions from M1 are more homogeneous than M2 and equally skewed as M3. In (c), M2’s predictions are more heterogeneous than M3’s.
(a) M1 vs. M2
(b) M1 vs. M3
(c) M2 vs. M3
Figure A9: IWI predictions in overlapping locations in Rwanda (M3: CBw). Left: Correlation between the predictions of two models. Right: Overlapping locations. Colors indicate differences in predicted values per model pair. M2 generally underestimates wealth compared to M1 (blue in (a)) and M3 (red in (c)). Predictions from M1 are more homogeneous than M2 and equally skewed as M3. In (c), M2’s predictions are more heterogeneous than M3’s.
(a) M1 vs. M2
(b) M1 vs. M3
(c) M2 vs. M3
Figure A10: IWI predictions in overlapping locations in Gabon (M3: CNNa+CB). Left: Correlation between the predictions of two models. Right: Overlapping locations. Colors indicate differences in predicted values per model pair. M2 generally underestimates wealth compared to M1 (blue in (a)) and M3 (red in (c)). Predictions from M1 are more homogeneous than M2 and M3. In (c), M2’s predictions are more heterogeneous than M3’s.

### v3

@@ 0.A.2 Performance on test sets
Table A1: Reported validation R2 scores for each approach. M1 presents validation results for single-country training in Table 1 [17]. M3 reports validation results on census data, as shown in Fig. S4 of the supplementary information [2]. We retrain the M2 models for each country using a subset of the original features reported in [8], and report the performance (R2) of the best model, indicated by the lowest root-mean-square error, RMSE (see details in Section 2.2).
Country (Code) / R2 M1 M2 M3
Sierra Leone (SLE) 89.82 80.7 83.0
Liberia (LBR) 89.3 76.3 -
Uganda (UGA) 89.02 78.5 -
Rwanda (RWA) 87.41 67.5 82.0
South Africa (ZAF) 73.03 49.5 -
Gabon (GAB) - 81.9 -
Figure A11: Predicted wealth distributions. Every column compares the predicted wealth distributions between two models. M1 and M2 predict IWI scores, while M3 predicts RWI scores. For comparisons with M3, we apply a linear transformation from RWI to IWI using Equation 1, with M1 and M2 as baselines in the middle and rightmost columns, respectively. M3’s distribution (blue) aligns more with M2 (orange), as indicated by a lower Jensen-Shannon (JS) distance, suggesting similar skewness levels (Gini coefficient) between them.
0.A.3 Comparison against M3, Chi et al. (2022)
The Relative Wealth Index (RWI) provides relative wealth scores that allow for comparisons within a given country, offering insights into spatial wealth disparities. However, a key limitation of the RWI is its lack of comparability across countries and over time, as it reflects relative wealth only within the specific context of the data used either from survey data or to train ML models. Consequently, RWI scores are not directly comparable across different geographic regions or time periods due to variations in the underlying data and economic conditions. To address this limitation, some studies have proposed transformations of the RWI to create a more standardized and comparable index across countries and years. One such proposal is the Comparable Wealth Index (CWI) [25], which aims to standardize wealth measurement globally. The CWI adjusts RWI scores to account for cross-national differences, allowing for more meaningful comparisons across different socioeconomic contexts. For a detailed comparison of wealth scores derived from survey data, see [20].
Although conversion methods between wealth indicators have been proposed, they commonly rely on reference measures that are not easily available. For this reason here we decided to follow a simple approach to fairly compare different methods. First, we make the assumption that M3 (which uses RWI) correlates strongly with M1 (which uses IWI), as evidenced by the very high correlation between RWI and IWI on survey data (Pearson’s correlation: $0.93-0.99$ [27]). Under this assumption, we apply a linear transformation of RWI to IWI using Equation 1, enabling us to make relative comparisons between M1 and M3. Similarly, we apply the transformation when comparing M2 to M3 using M2 as a baseline. While this transformation is not ideal and may introduce some inaccuracies, it serves as a useful tool for comparing the shape of the predicted distributions. These results should therefore be interpreted with caution.
Figure A11 shows the comparison of predicted wealth distributions across models. We observe that the shape of M3’s distributions and its wealth inequality more closely align with M2, as indicated by a smaller Jensen-Shannon distance [19] and similar Gini coefficient values compared to M1, suggesting that both models (M3 and M2) produce similar distributions of wealth. One possible explanation for this behavior is the similarity in the models’ architectures and the features they use, see Table 1. Note that to preserve the integrity of the data provided by the three models, we chose to retain outliers in the distributions. This decision results in apparent mismatches between the minimum and maximum values across distributions in Figure A11. Removing these outliers would lead to a more favorable but unfair comparison of the altered distributions.